OpenStudy (fernandez555):

reward+fan please help! )): Given the trinomial 2x^2 + 4x + 4, predict the type of solutions. A. 2 rational solutions B. 1 rational solution C. 2 irrational solutions D. 2 complex solutions

OpenStudy (fernandez555):

@ajprincess @mathstudent55 @fiercesteak @amoodarya

OpenStudy (campbell_st):

use the discriminant

OpenStudy (amoodarya):

factor 2 turn into a(x-b)^2+c

OpenStudy (amoodarya):

2x^2+4x+4 =2 {x^2+2x+2} =2{(x+1)^2-1+2}

OpenStudy (campbell_st):

its b^2 -4ac if b^2 - 2ac > 0 2 unequal real roots, if its a perfect square, the 2 roots are rational b^2 - 4ac = 0 2 equal real roots b^2 - 4ac < 0 2 complex roots